(-1.4*223,000-26100)*(1-0.35)+26,100
=?
Answer #1

1.4 * 223,000 = -$312,200

(-1.4 * 223,000 - $26,100) = -$338,300

-$338,300 * (1 - 0.35) = - $219,895

- $219,895 + $26,100 = - $193,795

Answer is - $193,795